Richard Gaziano is a manager for Health Care, Inc. Health Care deducts Social Security, Medicare, and FIT (by percentage method) from his earnings. Assume a rate of 6.2% on $118,500 for Social Security and 1.45% for Medicare. Before this payroll, Richard is $1,000 below the maximum level for Social Security earnings. Richard is married, is paid weekly, and claims 2 exemptions.

What is Richard’s net pay for the week if he earns $1,100?
Answer #1

Richard Gaziano

Before this payroll, Richard is $1,000 below the maximum level for Social Security earnings. Richard is married is paid weekly, and claims 2 exemptions.

The Richard's net pay for the week if he earns $1,100

The Social security =6.2/100 *1000 = 62

The medical care=1.45/100 *1000 = 14.5

So he will earn a net pay of =62+14.5+1100 = $1176.50